OpenStudy (anonymous):

The cost of 3 notebooks and four pencils is $8.50. The cost of 5 notebooks and 8 pencils is $14.50. Determine the cost of one notebook and the cost of one pencil. Only in algaebraic solution.

OpenStudy (anonymous):

let notebook = x and pencil =y so 3x+4y=8.50 and 5x+8y=14.50 so now that you have formed your simultaneous equations, you can find x and y

OpenStudy (anonymous):

thanks but can u explain to me how to solve equations like this thank You

OpenStudy (anonymous):

alright, sure so first of all, we need to either make the x's the same or the y's the same like 3x and 5x should be a common number such as 15x

OpenStudy (anonymous):

or to make it easier, you can just make the y's the same by multiplying the first equation by 2 in order to 8y

OpenStudy (anonymous):

2(3x+4y=8.50) = 6x+8y=17 <---1st equation 5x+8y=14.5 <---- 2nd equation

OpenStudy (anonymous):

now that your y's are the same, you can cancel them out by subtracting, meaning 8y-8y = 0 now since you subtracted the y's you have to subtract everything else also

OpenStudy (anonymous):

meaning it would be 6x-5x=17-14.5 so x=2.5

OpenStudy (anonymous):

after this, you substitute this value of x into any one of the above equations to get your value of y : (I'm using the first equation) : 6(2.5) + 8y =17 so 15+8y =17 hence 8y = 2 and your y is 2/8 which is 1/4 or 0.25 in decimals

OpenStudy (anonymous):

Thank You soooo very much :D

OpenStudy (anonymous):

and to make sure that your answer is correct, you can substitute your calculated x and y values back into both equations to see if you get the right answers

OpenStudy (anonymous):

you're welcome :) i hope you understood it

OpenStudy (anonymous):

U just gained a fan :D